OpenStudy (anonymous):

Solve the initial value problem: dx/dt = 10x-x^2 , x(0) =1

OpenStudy (anonymous):

use separable variable method dx/ (10x-x^2)=dt integrate and find constant C

OpenStudy (anonymous):

i should say constant C(t)

OpenStudy (anonymous):

Before integrating I got \[dt = \frac{ 1 }{ 10x } + \frac{ 1 }{ 10(1-x) }\] Is that right?

OpenStudy (anonymous):

method of partial fractions or try to make it a perefect square, if thats easy for you, many methods are available,its your call :)

OpenStudy (anonymous):

Answer in the book is \[x(t) = \frac{ 10 }{ 1+9e ^{-10t} }\] I don't known how they got that. My answer is t(something) not x(something), and the ln's make 0 invalid so I'm completely lost.

OpenStudy (anonymous):

how did you intergrate 10x-x^2

OpenStudy (anonymous):

Partial fraction, see above.

OpenStudy (anonymous):

you should split that up as x(10-x), use partial fractions example 1/(10x-x^2)= (a/x)+(b/10-x)

OpenStudy (anonymous):

You see the post above where I have the result of my partial fraction right?

OpenStudy (anonymous):

no you do not, you taken out 10 as a common factor which is not the case,

OpenStudy (anonymous):

Ok how about this: \[dt = \frac{ 1 }{ x } + \frac{ 1 }{ 10-x }\] \[t = \ln|x| - \ln|10-x| +c\]

OpenStudy (anonymous):

No, wait, the 10 in the denom came as a result of 10A=1 from the partial fraction process not a common factoring error.

OpenStudy (anonymous):

b is also 1/10

OpenStudy (anonymous):

i checked your work, and values of a and b are fine

OpenStudy (anonymous):

k so now how do I use that to get the answer in the book posted earlier.

OpenStudy (anonymous):

what are the values of integrals?

OpenStudy (anonymous):

you are almost there, the book uses log properties like say ln (ab)= ln(a)+ln(b) and also ln(a)-ln(b)=ln(a/b)

OpenStudy (anonymous):

\[t = \frac{ \ln|x| }{ 10 } - \frac{ \ln|x-10| }{ 10 }\]

OpenStudy (anonymous):

remember you need to find value of constant c(t), put t=0 and find value of contant c(0), as you already know that x(0)=1

OpenStudy (anonymous):

So t = 0 and x = 1?

OpenStudy (anonymous):

tell me what do you understand by x(0)=1, ? x is a function of t, so when t=0, x(0)= 1, does that help you will need to use this concept to find value of constant C, without that your answer is incorrect, or couple of points will be docked off on your test : )

OpenStudy (anonymous):

i hope that makes sense : )

OpenStudy (anonymous):

So I get: \[c(t) = \frac{ \ln|9| }{ 10 }\] Is that correct?

OpenStudy (anonymous):

If I put that into the general soln, I get a big mess, and it looks nothing like the book answer :(

OpenStudy (anonymous):

And if x is a function of t, why did I get t being a function of x after integrating. Where is x(t) even coming from?

OpenStudy (anonymous):

Three hours and still on the first question, with tons of other HW to do, awesome.

OpenStudy (turingtest):

ok, you have it up to here:\[t = \frac{ \ln|x| }{ 10 } - \frac{ \ln|10-x| }{ 10 }\](you switched the -10 and the x for some reason, but since its absolute value I guess it doesn't matter. I would simplify this using\[\ln a-\ln b=\ln\frac ab\]\[t=\frac1{10}\ln|\frac x{10-x}|\]

OpenStudy (anonymous):

Yes, I got that far, but the answer is in terms of x's, and the book soln is in terms of t's...

OpenStudy (turingtest):

sorry forgot the thing we were looking for, the integration constant C\[t=\frac1{10}\ln|\frac x{10-x}|+C\]\[x(0)=1\]so plug in t=0 and x=1\[0=\frac1{10}\ln|\frac19|+C\implies C=\frac{\ln 9}{10}\]so the problem must just be the form of the answer, since you seem right about C

OpenStudy (anonymous):

I made my calc solve for x, it gave:\[x = \frac{ 10e ^{10t} }{ e ^{10t}+9 }\] Close, but not it..

OpenStudy (anonymous):

I get it if I divide top and bottom by 1/e^10t, is that a valid maneuver?

OpenStudy (turingtest):

\[t=\frac1{10}\ln|\frac x{10-x}|+\frac{\ln9}{10}=\frac1{10}\ln|\frac{9x}{10-x}|\]\[\exp(10t)=\exp\left(\ln|\frac{9x}{10-x}|\right)\]\[e^{10t}=\frac{9x}{10-x}\]\[10e^{10t}-e^{10t}x=9x\]yes it is!

OpenStudy (turingtest):

so that's the only problem? lol cuz I was gonna say, the answer is right

OpenStudy (turingtest):

\[t=\frac1{10}\ln|\frac x{10-x}|+\frac{\ln9}{10}=\frac1{10}\ln|\frac{9x}{10-x}|\]\[\exp(10t)=\exp\left(\ln|\frac{9x}{10-x}|\right)\]\[e^{10t}=\frac{9x}{10-x}\]\[10e^{10t}-e^{10t}x=9x\]\[9x+e^{10t}x=10e^{10t}\]\[x=\frac{10e^{10t}}{9+e^{10t}}=\frac{10}{9e^{-10t}+1}\]

OpenStudy (turingtest):

is that what you were looking for?

OpenStudy (anonymous):

Your terms in the bottom are backwards to the answer, but I'm sure due to some arbitrary math rule its equivalent. I guess it's solved, I'm not sure I know what I just did or what any of this means, but that's not your problem.

OpenStudy (turingtest):

but the 9 is next to the e^(-10) right?

OpenStudy (anonymous):

Yup.

OpenStudy (turingtest):

so it's just x=10/(1+9e^(-10t)) yes you should definitely realize that a+b=b+a, so those two are the same

OpenStudy (anonymous):

:) I was imaging a subtraction between them.

OpenStudy (turingtest):

ah, yeah algebra never gets to be very fun because of things like that ;)

OpenStudy (anonymous):

K, thanks for coming to the rescue, on the the next problem.

OpenStudy (turingtest):

Post it if you are having trouble I'll do what I can, but I am learning my own stuff too :)

OpenStudy (turingtest):

post separately of course ...

OpenStudy (anonymous):

Roger that.